# On the Onset of Coherent Phonon Motion in Peierls-Distorted Antimony by Attosecond Transient Absorption

### v2

@@ Introduction
The coupling of optical excitation to atomic forces and lattice deformation offers intriguing ways to reach non-equilibrium and meta-stable states, with interesting proposed applications including electronic1, 2 and magnetic switching3, 4, 5.
The use of coherent optical pulses allows these transformations to occur on ultrafast timescales and with potentially higher efficiency compared to incoherent thermal pathways 6. To optimally develop and engineer these non-thermal pathways towards out-of-equilibrium phases, precise understanding of the underlying mechanisms in which the optical excitation couples to phonon motion needs to be developed.
While coherent phonons can be created directly out of coupling to the electromagnetic radiation, e.g. from THz frequency photons or Raman excitation by higher frequencies, an indirect mechanism termed “displacive excitation of coherent phonons” (DECP) is known to be prominent 7 and was shown as a promising route to control phonon motion 8. In this mechanism the photon-driven electronic excitation changes the energetic landscape of the solid, leading the lattice to rearrange to a new energetic optimum position, launching an ultrafast coherent optical phonon motion.
@@
In their original work on DECP, Zeiger et al. 7 describe a simple model of this motion and successfully explain the observed temporal behavior of the NIR reflectivity due to atomic forces acting on the lattice following optical excitation. Early follow up work identified a deviation in the phase of the phonon motion in antimony (Sb) that could not be explained by this displacive model 9, showing limitations in the understanding of the creation of this motion. Recent experimental and theoretical work has identified additional forces acting on the lattice during the relaxation of optically excited carriers as a potential source of this phase 10, 11, i.e. the excited carriers exact a different force on the lattice directly following optical excitation compared to when they have reached a thermal distribution. The theoretical results implied that these additional forces could affect the efficiency of the conversion of optical excitation into atomic motion, which would be desired for the proposed applications. To address measurements of such a phase, extremely short timescales are required. Recent experiments using time-resolved spectroscopy with temporal resolutions on the order of a few femtoseconds (fs) of metals 12, semiconductors 13, 14, 15 and semimetallic graphite 16, 17, have shown that there can be significant observable differences between the creation of excited carriers by a pump pulse and the formation of a thermalized distribution through carrier-carrier and carrier-phonon scattering (carrier thermalization). In contrast in the derivation of the model equations, Zeiger et al. note that both the number of optically (i.e. non-thermal) excited electrons and the change of electron temperature lend themselves as the basis for their model and that they expected equal outcomes for both. This leads to the opportunity to explore whether carrier relaxation affects the coherent phonon motion.
In this work, we explore the behavior of the coherent phonon signal following displacive excitation by an NIR pulse of Peierls-distorted Sb using attosecond XUV transient absorption spectroscopy, which has been shown to be sensitive to coherent phonon motion through energetic shifts of the observed absorption 18, 19.
@@
This work significantly extends a previous study of the XUV transient reflectivity spectroscopy of the coherent phonon motion of Peierls-distorted Bi 19. The component analysis presented here gives equivalent results compared to a spectral modelling fitting as introduced in previous XUV transient spectroscopy studies 19, 13, 14, 20. The increased experimental sensitivity and time-delay sampling in this study allows the detailed exploration of the early-timescale dynamics, elucidating the role of carrier relaxation on the phonon motion.
@@ Results
In Sb, the Peierls distortion leads to a displacement of neighboring hexagonal planes along the c lattice coordinate, with the so-called Peierls parameter at 298 K (whereas would correspond to equidistant spacing of hexagonal planes and a cubic lattice symmetry) 21. The distortion leads to an opening of a pseudo-gap at the Fermi energy () and the Sb semimetallicity, which exhibits a narrow density of states (DOS) around - characteristic for the group V semimetals 22. In addition to the original DECP studies on Bi and Sb, coherent phonon motion is well studied by transient reflectivity and absorption in the NIR to visible 7, 9, 23, 24, 25 and XUV spectral regions 19, as well as photoelectron spectroscopy 26 and time-resolved diffraction 27, 28, 29. Recent theoretical studies using density-functional theory 30, 31, 11 (DFT) and molecular dynamics 32 (MD) simulations have successfully treated the excitation of the coherent phonon motion due to a change of the minimum coordinate of the potential energy surface as a function of electron temperature or number of excited electrons.
A band-structure diagram and (projected) density of states plot (DOS) of Sb at equilibrium, obtained via DFT (see Appendix B), can be seen in Fig. 1a,b. Typical for the semimetallic properties, the density of states is minimal around the Fermi level with small pockets below the Fermi level around the L-point. The bands mostly exhibit -orbital character, indicating excellent spectroscopic visibility from the core levels. For the spectroscopic experiments, polycrystalline thin film samples of Sb of 32 nm thickness on Si3N4 membranes are prepared (see Appendix A). The laser pulses used for the experiments are 4 fs duration near infrared pulses from a Ti:sapphire laser for excitation and a short pulse train of attosecond pulses in the extreme ultraviolet for the probe, produced by the process of high harmonic generation (see Appendix C).
@@
In Fig. 1c, the static absorbance of Sb in the region between 29 eV and 39 eV is shown. Dominating are two peaks at 32.6 eV and 33.9 eV, corresponding to transitions of electrons from the 4d5/2 and 4d3/2 (semi-)core levels of Sb in the conduction bands (CB). Towards lower photon energies a falling background absorption can be observed that is associated with high-energy valence transitions (i.e. not core-level transitions). The absorption onset leads into a plateau region of absorption with additional peaks at 35 eV and 36 eV. The structures agree well with previous experiments on the XUV absorption spectrum of Sb 33, from which we accordingly assign the core-to-Fermi-level transition energies at 32.1 eV and 33.4 eV for the 4d5/2 and 4d3/2 levels, respectively. Using spin-orbit decomposition (see Appendix E.2), the absorption from the core-level is reconstructed. By comparing to the DOS (cf. Fig. 1b), we can associate these peaks with high-symmetry points in the conduction bands, such as around the point.
@@
Next, the long-time-range transient absorption of Sb is taken. Spectra are recorded for time delays between –0.2 ps and 4.0 ps in 6.7 fs steps, with an estimated incident fluence of 4.4 mJ/cm2. The fluence was chosen to obtain a high signal-to-noise ratio while staying sufficiently below sample degradation. The Gaussian instrument response function (IRF), dominated by the duration of the pump pulse, is determined to be fs (FWHM, see Appendix D). A small residual signal is observed at delay times where the NIR pulse arrives long after the XUV pulse. From our previous knowledge of measurements on thin film samples, this is typically due to a laser-induced heating. This small contribution is subtracted equally from all spectra. The resulting delay-dependent change in absorbance is shown in Fig. 1f (for a complete description of data acquisition and analysis, see Appendix E). At delay settings where the XUV probe precedes the NIR pump, no significant change in absorbance is visible. At the zero-time delay, a strong change in the signal is observed. An increase in absorbance occurs at energies below the onset of the 4d to CB absorbance of the static spectrum and a decrease in absorbance at energies above. The feature is repeated for both the spin-orbit split 4d5/2 and 4d3/2 orbitals. Furthermore, changes in absorbance are observed both within the energy range around accessible by the broad band NIR pump (e.g. roughly below 36 eV, see shaded area in Fig. 1(a)), and above. This shows that the observed dynamics are both due to a change in allowed core-level transitions following a pump-induced change in band-occupations (state opening/blocking) and overall changes in the electronic band structure. At early timescales the increased absorption features extend to lower energies compared to later timescales, indicating a rapid rearrangement of the carriers within the first hundreds of femtoseconds after excitation. Additionally, the induced changes slowly decay towards longer time delays. Most strikingly, as observed in other investigations 18, 19, 34, 35, the change in absorbance exhibits strong oscillatory features at positive delay times. These oscillations appear to modulate the observed changes in absorption, which quickly decay towards greater time delays, up to around 1.0 to 1.5 ps, i.e. faster than the overall decay of the pump-induced changes in absorption. These can be associated with the coherent phonon motion.
@@
To allow further insight into the observed dynamics and to reduce the complexity of the observed signal, a component analysis approach via singular value decomposition (SVD) is chosen. A short introduction into the method is given in Appendix E.3. The SVD yields “global” spectral and temporal components from the experimental data, i.e. without the need to further select certain temporal or spectral ranges and without fitting of an underlying physical model. To perform the SVD in the basis of the static absorbance, the static absorbance is added back to the measured change in absorbance. The leading component, encoding the static absorbance, is then omitted from the following analysis. The resulting spectral components (Fig. 1d,e) and temporal components (Fig. 1g) for the next two strongest components are resolved (a discussion of the residual components is presented in the SI): Analyzing the change in absorbance of the spectral components, we note that the first component (Fig. 1d) shows a strong, asymmetric lineshape around the Fermi-energy of the core-level transitions (32.1 eV and 33.4 eV). This is consistent with a broadening of the Fermi-Dirac distribution due to an increase in carrier temperature. Indeed, fitting a Fermi-Dirac distribution (convoluted with the DOS) for an increase in peak carrier temperature of 2000 K to both spin-orbit core-to-Fermi transitions yields good agreement with the obtained distribution and the thermal energy of the absorbed pump pulse. On the other hand, the second component (Fig. 1e) shows a similar asymmetric lineshape around the energy of the absorption peaks. This is consistent with an overall shift of the absorption spectrum due to movement of the lattice, which has been found to be the predominant effect of the coherent phonon motion 19. This can be confirmed by comparing the component to the change in absorbance that a shifted static absorption would yield. Good agreement is found for a shift of the static of -28 meV at the largest excursion. Therefore, the first component is associated with the change in carrier temperature and the second component with the dynamic of the lattice motion. Small deviations of this can be observed: In case of the lattice signal, while recent DFT calculations for Bi 19 have shown that the change of the Peierl’s coordinate leads to an energy shift of the absorption, it also can lead to rearranging of the absorbance due to changes in the electronic structure and transition dipole moments. The component analysis could potentially be better suited to represent these changes compared to a fitted modelling approach, which considers only changes from a shifted absorption. In the carrier signal, differences between the hot Fermi-Dirac distribution and the observed spectrum could be due to different transition strengths between populations in different bands. At higher energies, a broad feature around 36.5 eV is observed in the carrier component. These photon energies are above the resonant one-photon excitation window of the pump pulse from the VB. It is unlikely that they originate from a non-linear or multiphoton excitation process, as an occupation of states high in the CB would be expected to yield a negative change of absorbance, compared to the observed increase. A possible origin of this feature is a change of band structure in the higher valence bands due to photodoping by the excited carriers 36.
@@
The observation seems to indicate that the component analysis follows a two temperature model (TTM), where a laser-induced heating of the carrier temperature exchanges energy with the lattice subsystem. This is furthermore supported by the temporal dynamics: We observe a strong rise and subsequent exponential decay following zero-delay in the first component (blue dots in Fig. 1g) and a strong, exponentially damped oscillatory modulation and a slow rise in the second component (orange dots). The fast rise and exponential decay of the first component is characteristic for an effect of a change in carrier temperature, while the oscillatory behavior and the slow rise of the second component can be associated with a change of lattice structure due to the phonon motion and the heating of the lattice. The carrier temperature signal appears to be modulated by the phonon oscillation, and the maxima in the carrier temperature coincide with minima in the lattice signal and vice-versa. A similar effect had been previously theoretically predicted 30 and was observed in the XUV transient reflectivity spectroscopy of bismuth 37, where the result was associated with a modulation of the signal due to the back-coupling interaction of the lattice shift by the coherent phonon motion on the specific heat capacity of the carriers and therefore their temperature.
@@
Notably, however, and in conflict with a two-temperature model, the signal associated with a change in carrier temperature does not immediately reach its maximum within the instrument response function (IRF), but appears delayed (see shaded area in Fig. 1g). While a delayed build-up dynamic, as observed in the carrier-signal here, has been observed in reflectivity measurements in Bi and Sb before, it has been largely associated with a delayed response of the electronic system, but has found little discussion. Similar effects have been observed for XUV transient absorption spectroscopy in semiconductors 13, 14, 15, where they were related to relaxation of optically excited carriers to a thermodynamic equilibrium via intra-valley carrier-carrier scattering or inter-valley carrier-phonon scattering. On the other hand, Giret et al. 30 have discussed effects of carrier diffusion along the layer thickness in X-ray diffraction experiments 27. Extending a two-temperature model to include one-dimensional carrier diffusion to explain the delay, we are unable to find a sufficient diffusion effect on the carrier temperature, even when varying reported diffusivity values for Sb over orders of magnitudes. This is also supported by evaluating the residual components, which show a large signal in the first 200 fs following photoexcitation, implicating that the carrier occupation does not follow a thermal distribution at these early timescales (Residuals are shown in Appendix F and the carrier occupation following photoexcitation is discussed in more detail below). This shows, that the TTM is insufficient to describe the carriers during the first few femtoseconds following optical excitation, during which the carriers occupy highly excited states that do not follow a thermal distribution and are not in thermodynamic equilibrium with each other.
While more complex modelling based on time-dependent Boltzmann equations would be able to more accurately describe the band-dependent occupations during thermalization, we find that an extended two temperature model with a classical treatment of the phonon motion that treats the non-thermal carriers as an effective population (see Appendix G) sufficiently covers the observations and allows for a meaningful quantization of the observed dynamics within the component analysis. Within this model, the non-thermal carriers relax into the thermal pool of carriers and increase the effective temperature of the thermalized carriers. The best-fit result of a non-linear least-squares fit of the model to the carrier temperature and lattice dynamic is shown in Fig. 1g (purple and red solid lines). All obtained values are shown in the appendix (Table 1) and the time-dependent properties are discussed below.
@@
Compared to previous results of the Sb coherent phonon mode, measured by transient reflectivity studies in the NIR spectrum (obtaining 4.5 THz) 7, a smaller value of the A1g phonon frequency of 4.190.01 THz is observed from the fitting for these particular thin film samples. Similarly, a longer lifetime for the carrier recombination (2.950.04 ps) and a much shorter lifetime of the coherent phonon motion (0.960.04 ps) is obtained, compared to reference data (1.67 ps and 2.90 ps, respectively) 7. There are several possible explanations for these differences. First, as the observation of a residual heat signal shows, the very-thin samples are heated by the laser excitation to temperatures above room temperature. It was shown that higher temperatures lead to a softening of the phonon frequency, as well as a faster decay of the coherent phonon in Sb 24, as well as in Bi 38 and Ti2O3 39. However, in the case of Ti2O3 the increase in temperature was also associated with a fast recombination of carriers, which is opposite to the elongated lifetime of carriers observed here. Secondly, a similar effect on the softening and phonon lifetime has been observed to be dependent on the pump fluence for Sb 40, Bi 41 and Te 42. Lastly, the (polycrystalline) thin films are known to influence the coherent phonon mode as well, due to their higher defect density compared to single crystals and also due to confinement within the films. However, the latter effects would lead to a shift in frequency for the A1g mode in the Raman spectrum as well, which was found to be close to literature reference values (4.4 THz in the Raman here vs. 4.5 THz reported 7). Extrapolating changes in the phonon decay rate in Sb that have been observed for temperatures below 300 K to higher temperatures below the melting point of Sb would not yield the observed results 24. Comparison to measured fluence dependencies of the phonon frequency and decay rate by NIR single shot transient absorption spectroscopy on Sb thin films 40 yields good agreement with the values obtained here.
@@
The phase of the phonon oscillation from the fitting procedure (-0.09±0.02 rad) clearly differs from the pure DECP model, which only predicts a phase-effect from the damping (which we find accounts to rad) 7. This indicates a delayed reaction of the phonon motion. Reported values on the A1g phase in the literature are varying, ranging from -0.020.02 rad 7 to -0.130.01 rad 9 and warrant a closer inspection later.
@@
We do find good agreement between the modeled delayed rise of the carrier temperature due to carrier thermalization to the experimental observation, however it is necessary to include a scaling factor that limits the effect of the thermalization to achieve a good fit. The observed relaxation timescale () is furthermore comparable to previous observations in NIR reflectivity 23. Here we use the word thermalization to refer to the possibility of several processes, carrier-carrier scattering and carrier-phonon scattering, reaching a hot Fermi Dirac distribution and these can be carrier-energy dependent. While relaxations via carrier-carrier scattering have reported lifetimes of a few femtoseconds, relaxation via carrier-phonon scattering has been reported to occur on tens to hundreds of femtoseconds, making it the likely dominating observed mechanism. The apparent effect of this scaling factor could be due to two different effects: Either part of the excited carrier population relaxes into the thermalized carrier distribution within the duration of the pump pulse, while other carriers thermalize on measurably longer timescales, or a part of the excited carrier population already occupies the same states that are within their thermal distribution. Recent relaxation rate calculations by O’Mahony 40 show a long lifetime for carriers at or below , comparable to the result of the fit (), but predict fast relaxation ( fs) for carrier energies eV above or below . This is due to the larger density of states in those regions, leading to a much faster carrier relaxation than in the narrow regions closer to .

### v3

@@ Introduction
The coupling of optical excitation to atomic forces and lattice deformation offers intriguing ways to reach non-equilibrium and meta-stable states, with interesting proposed applications including electronic[1, 2] and magnetic switching[3, 4, 5].
The use of coherent optical pulses allows these transformations to occur on ultrafast timescales and with potentially higher efficiency compared to incoherent thermal pathways [6]. To optimally develop and engineer these non-thermal pathways towards out-of-equilibrium phases, precise understanding of the underlying mechanisms in which the optical excitation couples to phonon motion needs to be developed.
While coherent phonons can be created directly out of coupling to the electromagnetic radiation, e.g. from THz frequency photons or Raman excitation by higher frequencies, an indirect mechanism termed “displacive excitation of coherent phonons” (DECP) is known to be prominent [7] and was shown as a promising route to control phonon motion [8]. In this mechanism the photon-driven electronic excitation changes the energetic landscape of the solid, leading the lattice to rearrange to a new energetic optimum position, launching an ultrafast coherent optical phonon motion.
@@
In their original work on DECP, Zeiger et al. [7] describe a simple model of this motion and successfully explain the observed temporal behavior of the NIR reflectivity due to atomic forces acting on the lattice following optical excitation. Early follow up work identified a deviation in the phase of the phonon motion in antimony (Sb) that could not be explained by this displacive model [9], showing limitations in the understanding of the creation of this motion. Recent experimental and theoretical work has identified additional forces acting on the lattice during the relaxation of optically excited carriers as a potential source of this phase [10, 11], i.e. the excited carriers exact a different force on the lattice directly following optical excitation compared to when they have reached a thermal distribution. The theoretical results implied that these additional forces could affect the efficiency of the conversion of optical excitation into atomic motion, which would be desired for the proposed applications. To address measurements of such a phase, extremely short timescales are required. Recent experiments using time-resolved spectroscopy with temporal resolutions on the order of a few femtoseconds (fs) of metals [12], semiconductors [13, 14, 15] and semimetallic graphite [16, 17], have shown that there can be significant observable differences between the creation of excited carriers by a pump pulse and the formation of a thermalized distribution through carrier-carrier and carrier-phonon scattering (carrier thermalization). In contrast in the derivation of the model equations, Zeiger et al. note that both the number of optically (i.e. non-thermal) excited electrons and the change of electron temperature lend themselves as the basis for their model and that they expected equal outcomes for both. This leads to the opportunity to explore whether carrier relaxation affects the coherent phonon motion. In this work, we explore the behavior of the coherent phonon signal following displacive excitation by an NIR pulse of Peierls-distorted Sb using attosecond XUV transient absorption spectroscopy, which has been shown to be sensitive to coherent phonon motion through energetic shifts of the observed absorption [18, 19].
While it is well known that the relaxation of carriers in solids are generally many-body functions, i.e. they not only depend on their momentum and band index or energy, but also that of all other carriers, several approximations can often be made to simplify their description [20]. While a two-temperature model simply assumes the distribution of carriers to always represent a thermal (i.e. Fermi-Dirac) distribution and interacts thermally with the phonon population, this assumption necessarily breaks down at early-enough time-scales after photo-excitation where carriers might be energetically distributed far out of their thermal equilibrium. A first step to regain the complexity of the underlying dynamics is to extend the TTM by introducing a population of non-thermal carriers that decays with a constant decay rate, which we will call the extended TTM here. Further inclusion of the energy and momentum dependent relaxation rates, state- and population density leads to the popular Boltzmann equation model.
As we will see in this work, while the TTM well describes the transient XUV absorption dynamics of Sb, at early timescales it indeed brakes down and requires the introduction of the extended TTM. We will further see, that while the spectrally dependent carrier relaxation rates are not directly observable within the carrier population signal within the signal-to-noise rate of our experiment, they still observably influence the phonon motion.
@@
This work significantly extends a previous study of the XUV transient reflectivity spectroscopy of the coherent phonon motion of Peierls-distorted Bi [19]. The component analysis presented here gives equivalent results compared to a spectral modelling fitting as introduced in previous XUV transient spectroscopy studies [19, 13, 14, 21]. The increased experimental sensitivity and time-delay sampling in this study allows the detailed exploration of the early-timescale dynamics, elucidating the role of photoexcitation and carrier relaxation on the phonon motion and the Peierls distortion.
@@ Results
In Sb, the Peierls distortion leads to a displacement of neighboring hexagonal planes along the c lattice coordinate, with the so-called Peierls parameter at 298 K (whereas would correspond to equidistant spacing of hexagonal planes and a cubic lattice symmetry) [22]. The distortion leads to an opening of a pseudo-gap at the Fermi energy () and the Sb semimetallicity, which exhibits a narrow density of states (DOS) around - characteristic for the group V semimetals [23]. In addition to the original DECP studies on Bi and Sb, coherent phonon motion is well studied by transient reflectivity and absorption in the NIR to visible [7, 9, 24, 25, 26] and XUV spectral regions [19], as well as photoelectron spectroscopy [27] and time-resolved diffraction [28, 29, 30]. Recent theoretical studies using density-functional theory [31, 32, 11] (DFT) and molecular dynamics [33] (MD) simulations have successfully treated the excitation of the coherent phonon motion due to a change of the minimum coordinate of the potential energy surface as a function of electron temperature or number of excited electrons.
A band-structure diagram and (projected) density of states plot (DOS) of Sb at equilibrium, obtained via DFT (see Appendix B), can be seen in Fig. 1e. Typical for the semimetallic properties, the density of states is minimal around the Fermi level with small pockets below the Fermi level around the L-point. The bands mostly exhibit -orbital character, indicating excellent spectroscopic visibility from the core levels. For the spectroscopic experiments, polycrystalline thin film samples of Sb of 32 nm thickness on Si3N4 membranes are prepared (see Appendix A). The laser pulses used for the experiments are 4 fs duration near infrared pulses from a Ti:sapphire laser for excitation (see Fig. 1b) and a short pulse train of attosecond pulses in the extreme ultraviolet for the probe (see Fig. 1a), produced by the process of high harmonic generation (see Appendix C).
@@
In the inset of Fig. 1c, the static absorbance of Sb in the region between 29 eV and 39 eV is shown. Dominating are two peaks at 32.6 eV and 33.9 eV, corresponding to transitions of electrons from the 4d5/2 and 4d3/2 (semi-)core levels of Sb in the conduction bands (CB). Towards lower photon energies a falling background absorption can be observed that is associated with high-energy valence transitions (i.e. not core-level transitions). The absorption onset leads into a plateau region of absorption with additional peaks at 35 eV and 36 eV. The structures agree well with previous experiments on the XUV absorption spectrum of Sb [34], from which we accordingly assign the core-to-Fermi-level transition energies at 32.1 eV and 33.4 eV for the 4d5/2 and 4d3/2 levels, respectively. Using spin-orbit decomposition (see Appendix E.2), the absorption from the core-level is reconstructed (see Fig. 2a). By comparing to the DOS (cf. Fig. 1e), we can associate these peaks with high-symmetry points in the conduction bands, such as around the point.
@@
Next, the long-time-range transient absorption of Sb is taken. Spectra are recorded for time delays between –0.2 ps and 4.0 ps in 6.7 fs steps, with an estimated incident fluence of 4.4 mJ/cm2, leading to an estimated photoexcited carrier density of cm, or approximately 1% of valence electrons. The fluence was chosen to obtain a high signal-to-noise ratio while staying sufficiently below sample degradation. The Gaussian instrument response function (IRF), dominated by the duration of the pump pulse, is determined to be fs (FWHM, see Appendix D). A small residual signal is observed at delay times where the NIR pulse arrives long after the XUV pulse. From our previous knowledge of measurements on thin film samples, this is typically due to a laser-induced heating. This small contribution is subtracted equally from all spectra. The resulting delay-dependent change in absorbance is shown in Fig. 1c (for a complete description of data acquisition and analysis, see Appendix E). At delay settings where the XUV probe precedes the NIR pump, no significant change in absorbance is visible. At the zero-time delay, a strong change in the signal is observed. An increase in absorbance occurs at energies below the onset of the 4d to CB absorbance of the static spectrum and a decrease in absorbance at energies above. The feature is repeated for both the spin-orbit split 4d5/2 and 4d3/2 orbitals. Furthermore, changes in absorbance are observed both within the energy range around accessible by the broad band NIR pump (e.g. roughly below 36 eV, see shaded area in Fig. 1(e) right), and above. This shows that the observed dynamics are both due to a change in allowed core-level transitions following a pump-induced change in band-occupations (state opening/blocking) and overall changes in the electronic band structure. At early timescales the increased absorption features extend to lower energies compared to later timescales, indicating a rapid rearrangement of the carriers within the first hundreds of femtoseconds after excitation. Additionally, the induced changes slowly decay towards longer time delays. Most strikingly, as observed in other investigations [18, 19, 35, 36], the change in absorbance exhibits strong oscillatory features at positive delay times. These oscillations appear to modulate the observed changes in absorption, which quickly decay towards greater time delays, up to around 1.0 to 1.5 ps, i.e. faster than the overall decay of the pump-induced changes in absorption. These can be associated with the coherent phonon motion.
@@
To allow further insight into the observed dynamics and to reduce the complexity of the observed signal, a component analysis approach via singular value decomposition (SVD) is chosen. A short introduction into the method is given in Appendix E.3. The SVD yields “global” spectral and temporal components from the experimental data, i.e. without the need to further select certain temporal or spectral ranges and without fitting of an underlying physical model. Here this global analysis is performed over the full spectral and delay range shown in Fig. 1c. To perform the SVD in the basis of the static absorbance, the static absorbance is added back to the measured change in absorbance. The leading component, encoding the static absorbance, is then omitted from the following analysis. The resulting spectral components (Fig. 2b,c) and temporal components (Fig. 2d) for the next two strongest components are resolved (a discussion of the residual components is presented in the SI): Analyzing the change in absorbance of the spectral components, we note that the first component (Fig. 2b) shows a strong, asymmetric lineshape around the Fermi-energy of the core-level transitions (32.1 eV and 33.4 eV). This is consistent with a broadening of the Fermi-Dirac distribution due to an increase in carrier temperature. Indeed, fitting a Fermi-Dirac distribution (convoluted with the DOS) for an increase in peak carrier temperature of 2000 K to both spin-orbit core-to-Fermi transitions yields good agreement with the obtained distribution and the thermal energy of the absorbed pump pulse. On the other hand, the second component (Fig. 2c) shows a similar asymmetric lineshape around the energy of the absorption peaks. This is consistent with an overall shift of the absorption spectrum due to movement of the lattice, which has been found to be the predominant effect of the coherent phonon motion [19]. This can be confirmed by comparing the component to the change in absorbance that a shifted static absorption would yield. Good agreement is found for a shift of the static of -28 meV at the largest excursion. Therefore, the first component is associated with the change in carrier temperature and the second component with the dynamic of the lattice motion. Small deviations of this can be observed: In case of the lattice signal, while recent DFT calculations for Bi [19] have shown that the change of the Peierls coordinate leads to an energy shift of the absorption. This is caused by a closing of the negative band gap and a renormalization of the conduction bands closer to the Fermi level, leading to an overall red-shift of the core-to-CB absorption. At the same time, the effect is counter-acted by the reduced Coulomb repulsion between the ions, which leads to a higher binding energy of the core-levels, which leads to a (smaller) blue-shift of the core-to-CB transition [19]. While the change in absorption can be approximated well by comparison the shifted static absorption, the component analysis applied here could potentially be better suited to represent the more detailed changes due to renormalization of the bands compared to a fitted modelling approach, which considers only changes from a shifted absorption. While the energy of the observed red-shift of the absorption due to the lattice motion is similar to the quasi-particle energy associated with the phonon oscillation, the two are (within first order) not connected, since the red-shift is dependent on the oscillation amplitude and not the frequency. In the carrier signal, differences between the hot Fermi-Dirac distribution and the observed spectrum could be due to different transition strengths between populations in different bands. At higher energies, a broad feature around 36.5 eV is observed in the carrier component. These photon energies are above the resonant one-photon excitation window of the pump pulse from the VB. It is unlikely that they originate from a non-linear or multiphoton excitation process, as an occupation of states high in the CB would be expected to yield a negative change of absorbance, compared to the observed increase. A possible origin of this feature is a change of band structure in the higher valence bands due to photodoping and many-body effects by the large number of excited carriers [37].
@@
The observation seems to indicate that the component analysis follows a two temperature model (TTM), where a laser-induced heating of the carrier temperature exchanges energy with the lattice subsystem. This is furthermore supported by the temporal dynamics: We observe a strong rise and subsequent exponential decay following zero-delay in the first component (teal dots in Fig. 2d) and a strong, exponentially damped oscillatory modulation and a slow rise in the second component (purple dots). The fast rise and exponential decay of the first component is characteristic for an effect of a change in carrier temperature, while the oscillatory behavior and the slow rise of the second component can be associated with a change of lattice structure due to the phonon motion and the heating of the lattice. The carrier temperature signal appears to be modulated by the phonon oscillation, and the maxima in the carrier temperature coincide with minima in the lattice signal and vice-versa. A similar effect had been previously theoretically predicted [31] and was observed in the XUV transient reflectivity spectroscopy of bismuth [38], where the result was associated with a modulation of the signal due to the back-coupling interaction of the lattice shift by the coherent phonon motion on the specific heat capacity of the carriers and therefore their temperature.
@@
Notably, however, and in conflict with a two-temperature model, the signal associated with a change in carrier temperature does not immediately reach its maximum within the instrument response function (IRF), but appears delayed (see shaded area in Fig. 2d). While a delayed build-up dynamic, as observed in the carrier-signal here, has been observed in reflectivity measurements in Bi and Sb before, it has been largely associated with a delayed response of the electronic system, but has found little discussion. Similar effects have been observed for XUV transient absorption spectroscopy in semiconductors [13, 14, 15], where they were related to relaxation of optically excited carriers to a thermodynamic equilibrium via intra-valley carrier-carrier scattering or inter-valley carrier-phonon scattering. On the other hand, Giret et al. [31] have discussed effects of carrier diffusion along the layer thickness in X-ray diffraction experiments [28]. Extending a two-temperature model to include one-dimensional carrier diffusion to explain the delay, we are unable to find a sufficient diffusion effect on the carrier temperature, even when varying reported diffusivity values for Sb over orders of magnitudes. This is also supported by evaluating the residual components, which show a large signal in the first 200 fs following photoexcitation, implicating that the carrier occupation does not follow a thermal distribution at these early timescales (Residuals are shown in Appendix F and the carrier occupation following photoexcitation is discussed in more detail below). This shows, that the TTM is insufficient to describe the carriers during the first few femtoseconds following optical excitation, during which the carriers occupy highly excited states that do not follow a thermal distribution and are not in thermodynamic equilibrium with each other.
While more complex modelling based on time-dependent Boltzmann equations would be able to more accurately describe the band-dependent occupations during thermalization, we find that an extended two temperature model with a classical treatment of the phonon motion that treats the non-thermal carriers as an effective population (see Appendix G) sufficiently covers the observations and allows for a meaningful quantization of the observed dynamics within the component analysis. Within this model, the non-thermal carriers relax into the thermal pool of carriers and increase the effective temperature of the thermalized carriers. The best-fit result of a non-linear least-squares fit of the model to the carrier temperature and lattice dynamic is shown in Fig. 2d (black solid lines). All obtained values are shown in the appendix (Table 1) and the time-dependent properties are discussed below.
@@
Compared to previous results of the Sb coherent phonon mode, measured by transient reflectivity studies in the NIR spectrum (obtaining 4.5 THz) [7], a smaller value of the A1g phonon frequency of 4.190.01 THz is observed from the fitting for these particular thin film samples. Similarly, a longer lifetime for the carrier recombination (2.950.04 ps) and a much shorter lifetime of the coherent phonon motion (0.960.04 ps) is obtained, compared to reference data (1.67 ps and 2.90 ps, respectively) [7]. There are several possible explanations for these differences. First, as the observation of a residual heat signal shows, the very-thin samples are heated by the laser excitation to temperatures above room temperature. It was shown that higher temperatures lead to a softening of the phonon frequency, as well as a faster decay of the coherent phonon in Sb [25], as well as in Bi [39] and Ti2O3 [40]. However, in the case of Ti2O3 the increase in temperature was also associated with a fast recombination of carriers, which is opposite to the elongated lifetime of carriers observed here. Secondly, a similar effect on the softening and phonon lifetime has been observed to be dependent on the pump fluence for Sb [41], Bi [42] and Te [43]. Lastly, the (polycrystalline) thin films are known to influence both the lifetimes of photoexcited carriers as well as the coherent phonon mode, due to their higher defect density compared to single crystals and also due to confinement within the films. While we do not observe trapping of photoexcited carriers due to defect states in the experiment directly, the observed timescales of the carrier dynamics are likely to be affected by the expected large defect density of the thin films. A direct comparison between a single crystalline sample and a thin-film polycrystalline sample could further elucidate this connection. For the coherent phonon frequency, effects due to defect density and confinement would lead to a shift in frequency for the A1g mode in the Raman spectrum as well, which was found to be close to literature reference values (4.4 THz in the Raman spectrum here vs. 4.5 THz reported [7]). Extrapolating changes in the phonon decay rate in Sb that have been observed for temperatures below 300 K to higher temperatures below the melting point of Sb would not yield the observed results [25]. Comparison to measured fluence dependencies of the phonon frequency and decay rate by NIR single shot transient absorption spectroscopy on Sb thin films [41] yields good agreement with the values obtained here.
@@
The phase of the phonon oscillation from the fitting procedure (-0.09±0.02 rad) clearly differs from the pure DECP model, which only predicts a phase-effect from the damping (which we find accounts to rad) [7]. This indicates a delayed reaction of the phonon motion. Reported values on the A1g phase in the literature are varying, ranging from -0.020.02 rad [7] to -0.130.01 rad [9] and warrant a closer inspection later.
@@
We do find good agreement between the modeled delayed rise of the carrier temperature due to carrier thermalization to the experimental observation, however it is necessary to include an additional scaling factor that limits the effect of the thermalization to achieve a good fit. This scaling factor describes that only about 60% of the signal associated with the hot carriers shows an observable relaxation timescale, while the remaining 40% seem to appear within the instrument response function (IRF). This implies either part of the excited carrier population relaxes into the thermalized carrier distribution within the duration of the pump pulse, while other carriers thermalize on measurably longer timescales, or a part of the excited carrier population already occupies the same states that are within their thermal distribution. Recent relaxation rate calculations by O’Mahony [41] show a long lifetime for carriers at or below , comparable to the result of the fit (), but predict fast relaxation ( fs) for carrier energies eV above or below . This is due to the larger density of states in those regions, leading to a much faster carrier relaxation than in the narrow regions closer to . The observed relaxation timescale () is furthermore comparable to previous observations in NIR reflectivity [24]. Here we use the word thermalization to refer to the possibility of several processes, carrier-carrier scattering and carrier-phonon scattering, reaching a hot Fermi Dirac distribution and these can be carrier-energy dependent. While relaxations via carrier-carrier scattering have reported lifetimes of a few femtoseconds, relaxation via carrier-phonon scattering has been reported to occur on tens to hundreds of femtoseconds, making it the likely dominating observed mechanism.